You drive to Sacramento which is 320 miles away in 5 hours. First, give your rate of traveling in miles per hour, then give the rate in minutes per mile.
Answer by Alan3354(69443)

You drive to Sacramento which is 320 miles away in 5 hours. First, give your rate of traveling in miles per hour,
-----
d = r*t
r = d/t
------
then give the rate in minutes per mile.
r from above = mi/hr
1/r = hr/mile
60/r = min/mile